OpenStudy (anonymous):

y + 4 x = 85 x + 2y = 13

OpenStudy (anonymous):

i dont know what you got but this is what i figured out... x=-13/77 Y=1053/77

OpenStudy (anonymous):

no

OpenStudy (anonymous):

First collect all y' s and x's put together